# VIPER: Visual Perception and Explainable Reasoning for Sequential Decision-Making

### v1

@@ 1.1 LLM and VLM for Sequential Decision-Making
When tackling long-horizon multimodal sequential decision-making tasks, embodied agents leverage key internal mechanisms such as perception and reasoning to achieve their objectives. Recently, LLMs and VLMs have demonstrated emergent capabilities in understanding complex situations, establishing them as essential architectures for embodied agents [ahn2022icanisay, shinn2023reflexion, yao2023react]. However, these models face challenges related to grounding and alignment with the physical world [mahowald2023dissociating]. A primary line of research focuses on leveraging LLMs for sequential decision-making through text, either by filtering irrelevant actions for a given task [yao2023react, shinn2023reflexion] or by learning from environmental feedback via reinforcement learning [carta2023grounding, tan2024true]. These methods have demonstrated impressive performance and generalization capabilities. However, they are primarily designed for textual environments, where all state information is transmitted by the environment through text, limiting their applicability to real-world embodied agents. Recently, research has increasingly focused on leveraging VLMs for sequential decision-making tasks in visual environments. For instance, EMMA [yang2024embodied] trained a VLM-based agent for sequential decision-making in visual environments by distilling the actions of a (SOTA) textual agents, such as Reflexion [shinn2023reflexion], within a parallel textual world. This approach represents an initial step toward utilizing multimodality for instruction generation. However, compared to our work, it remains dependent on a parallel textual environment for task learning, limiting its applicability to real-world scenarios. Closer to our work, RL4VLM [zhai2024fine] employs reinforcement learning to train VLM-based agents within visual environments, enabling them to learn to achieve tasks across diverse environments. However, its performance remains inferior to methods that rely solely on textual observation for reasoning and planning. To address these limitations, we propose a framework that leverages the reasoning capabilities of LLMs in the textual modality and applies them to real-world scenarios through the vision modality. Our approach VIPER creates a decomposition between perception and reasoning via an intermediate text modality: a VLM component processes visual input to generate textual descriptions of all elements present in a scene, and a fine-tuned LLM utilizes this description to generate actions that achieve the specified goal.
@@ 1.2 Text as Intermediate Modality
Recently, VLMs have demonstrated impressive performance on mainstream computer vision tasks such as object detection, visual question answering (VQA), and image captioning [ghosh2024exploring, zhang2024vision]. These tasks can be addressed using a single architecture and prompting, making VLMs a promising candidate for robotic applications. However, recent studies have shown that these architectures still underperform compared to their LLM counterparts on purely textual data. [wang2024pictureworththousandwords] demonstrated that VLMs often exhibit inferior reasoning capabilities compared to LLMs, highlighting the advantages of text-based reasoning. For example, [özdemir2024enhancingvisualquestionanswering] utilized a VLM for image captioning before performing a question-answering task to enhance image-related information and improve VQA performance, demonstrating the role of VLMs in extracting scene information. Similarly, [guran2024taskorientedroboticmanipulationvision] employed a VLM to extract bounding boxes and object information from a scene for robotic tasks, while [dang2025planningvisionlanguagemodelsuse] proposed Image-to-PDDL, where a VLM extracts a representation in Planning Domain Description Language (PDDL) [McDermott1998PDDLthePD] of the scene, which is then processed by an LLM to generate robotic actions, effectively leveraging the strengths of both architectures.
@@ 2.3 LLMs as Decision-Making Policies
We leverage the reasoning capabilities of LLMs to solve sequential decision-making tasks [carta2023grounding, shinn2023reflexion, tan2024true]. We follow the framework introduced in recent literature [carta2023grounding, tan2024true] to derive a policy (a probability distribution over possible actions) from an LLM, as shown in Figure 1(c). More precisely, given a prompt , the probability of an action sequence , denoted as , is obtained from the decoding probabilities of the tokens composing . Specifically, this probability is computed as:
The recent state-of-the-art RL4VLM agent [zhai2024fine] uses free-form generation of actions, including possible actions in the prompt. In contrast, the policy in Eq. 1 is directly constrained to the support of possible actions, improving performances on the downstream task (see Sec. 3).
In addition, while existing LLM agents methods [shinn2023reflexion, yao2023react, deps] receive prompts with textual observations directly from the environment, our method constructs the prompt using a textual description of the image observation generated by a VLM , along with the goal and the past actions.
@@ 2.4 Training Strategies
Recent works [carta2023grounding, aissi2024reinforcement] show that while LLMs possess common-sense knowledge, they lack grounding and are misaligned with physical environments [mahowald2023dissociating], leading to poor zero-shot performance in complex interactive decision-making tasks. To address this, following [carta2023grounding, tan2024true], we fine-tune the reasoning module (LLM) using a sequential strategy. First, we align it with a good reference policy using BC on a dataset of expert demonstrations, which is assumed to be available in the environment. Then, we refine its performance through RL via exploration and interaction with the environment. This BC+RL approach is well-established in the control community [lu2023imitation] and has demonstrated its relevance in our ablation study (see Section 3.3). While VIPER’s reasoning module operates on textual inputs, the textual description of the image observation is provided by a frozen VLM. Therefore, unlike EMMA [yang2024embodied], which trains its model by distilling a text expert policy and thus requires textual observations from the environment, our training relies solely on the goal and the image observations .
BC fine-tuning: We pre-train the LLM policy parametrized by a parameter , using a dataset of expert demonstrations of observation-action trajectories through BC. In our case, this expert dataset is created by selecting the successful trajectories of a rule-based expert (Sup. 5). The BC objective is to maximize the log-likelihood of expert actions with respect to the action distribution predicted by the LLM policy along the expert trajectory. Therefore, the BC loss is the negative log-likelihood, defined as:
@@
RL fine-tuning. We further enhance the policy through online interaction with the environment and fine-tune using Proximal Policy Optimization (PPO) [schulman2017proximal]. We learn a value function that approximates the true value for PPO. The policy is then optimized to minimize , as described in [schulman2017proximal].
@@ 2.5 Explaining Agent Decision
Furthermore, we perform weakly-supervised instance segmentation of important concepts in the image description used by the reasoning module by computing the integrated gradient [sundararajan2017axiomaticattributiondeepnetworks] of the chosen action with respect to each text token in the image description, and selecting the most significant ones. Following the same procedure, we compute the integrated gradient of the selected tokens, with respect to the image (see Sup. D for more details). This provides a rich tool for explainability, linking the performance of the perception and reasoning modules, and enabling to finely analyze failure cases (see Sec. 3.4).
@@ 3 Experimental Results
| Observation | Pick | Look | Clean | Heat | Cool | Pick2 | Avg | |
[…]
@@ 3.1 Experimental Setup
Environment. Our experiments are conducted on the ALFWorld benchmark [shridhar2021alfworldaligningtextembodied], a 3D environment that simulates six categories of household tasks with textual goals. ALFWorld has become a standard benchmark in the literature for evaluating models in long-horizon multimodal decision-making tasks, offering a realistic and interactive environment well suited for studying real-world embodied applications. This benchmark contains sets of training, in-distribution validation, and out-of-distribution (OOD) episodes that encapsulate specific layouts of the environment (room in the house, object and furniture placement, etc.). We follow the standard protocol and train VIPER on the training set of episodes, perform cross-validation on the in-distribution validation set, and report our results on the OOD test set. We evaluate the models using the mean success rate (higher is better, ) and the mean length of successful trajectories (lower is better, ). See Sup. 5 for more details about the environment.
Architecture and Training. VIPER uses Florence-2-220M [xiao2023florence] as VLM and Mistral-7B [jiang2023mistral7b] as LLM, with further discussions on model choices in Section 3.3. We fine-tune the LLM as described in Section 2.4 using low-rank adaptation (LoRA) with parameters and on the attention layers. For BC, we train the model until convergence and for PPO for approximately interaction steps with the environment. Both fine-tuning processes use a batch size of 2 and the Adam optimizer with a learning rate of . For each task, BC fine-tuning takes and PPO on 4 H100 GPUs.
Baselines. We compare VIPER with various SOTA models from the literature, categorized into two groups. The first group comprises models trained with supervision from a parallel textual environment, including Reflexion [shinn2023reflexion], DEPS [deps], ReAct [yao2023react], AutoGen [autogen] which only receive textual observation of the environment and EMMA [yang2024embodied] which uses both visual and textual observation during training but only access visual observation during inference. Like VIPER, the second group consists of models that exclusively rely on visual information from the environment, including zero-shot models (MiniGPT-4 [zhu2023minigpt4enhancingvisionlanguageunderstanding], InstructBLIP [dai2023instructblipgeneralpurposevisionlanguagemodels], Florence 2 [xiao2023florence], Idefics 2 [laurençon2024matters]) and a VLM fine-tuned through BC and RL, RL4VLM [zhai2024fine]. All methods of the first group accessing the textual observations benefit from richer scene information, including the number of objects, object IDs within the description, and the correspondence between object IDs and actions. In contrast, the models of the second group, including the zero-shot perception module in VIPER, cannot perfectly generate this level of detail, as this information is internal to the environment and not directly observable from images. Therefore, the models of the first group can be seen as oracle baselines.
@@ 3.2 Comparison with the state-of-the-art
We compare VIPER with the two baseline families on ALFWorld (see Table 1), evaluating both the mean success rate and the mean trajectory length required to achieve the goals. As observed, VIPER outperforms zero-shot visual-only methods, achieving an average 50pt improvement in success rate and an 8-step reduction in trajectory length across multiple tasks. It also surpasses RL4VLM [zhai2024fine], the SOTA method most closely related to our work that applies both BC and RL fine-tuning. This lower performance of RL4VLM [zhai2024fine] may be explained by the use of CoT reasoning with free generation for action prediction, which prevents it from being constrained to the action space. This leads to irrelevant predictions, as explained in Section 2.3. Another factor could be the absence of past action history in the agent’s prompt, as discussed in Sup. B.2. Besides, VIPER outperforms ReAct [yao2023react] and matches the performance of AutoGen [autogen] and DEPS [deps], which rely on textual observations from the environment. Moreover, VIPER approaches the performance of oracle methods such as EMMA [yang2024embodied] and Reflexion [shinn2023reflexion], achieving more than 70% success rate in the first five tasks and more than 50% in the Pick2 task. The performance gap between Pick2 and the other tasks results from the increased complexity of manipulating multiple objects. Text-based methods, such as EMMA and Reflexion, do not face this challenge, as they have access
@@
Comparisons between textual observations and VIPER-generated descriptions can be found in Sup. 5. Furthermore, VIPER achieves the shortest mean episode length compared to both the oracle and vision-only baselines, demonstrating that its successful trajectories are more efficient, taking more goal-directed actions and avoiding unnecessary steps.
3.3 Ablation study
Impact of BC and RL Training. Through an ablation study, we evaluate the impact of each training strategy defined in Section 2.4 – BC and RL – by comparing different variants of VIPER in Figure 2. Specifically, we report the performance of VIPER trained only with BC (VIPER BC) using rule-based expert demonstrations (see Sup. 5), only PPO (VIPER PPO), and the sequential approach where PPO follows BC (VIPER BC+PPO). Although using SOTA VLMs and LLMs, the zero-shot model performs poorly across all tasks, highlighting the complexity of the tasks and the necessity of fine-tuning. The results further show that while applying RL fine-tuning alone (VIPER PPO) is sufficient to learn simpler tasks (Pick and Look), using BC beforehand is essential to learn more complex ones (Clean, Heat Cool and Pick2). Furthermore, while VIPER BC already outperforms vision-only SOTA models, adding RL further improves performance, producing a boost of 3pt to 9pt depending on the task. This performance gain is larger when considering weaker reasoning modules (see Sup. C).

### v2

@@ Abstract
1 Introduction
The impressive capabilities of Large Language Models (LLMs) and Vision-Language Models (VLMs) are now well-demonstrated in various tasks, including respectively translation, summarization, and question answering [21, 31, 26] for Natural Language Processing (NLP) and object detection, image captioning, or visual question answering [7, 34] in computer vision.
These last years, a new line of research leverages recent advancements in LLMs and VLMs for the development of embodied AI agents [35, 30]. One example of task, tackled in this paper, concerns visual-based instruction-based planning, where an agent is tasked with solving a problem formulated through a textual prompt (e.g., “heat the apple and put it on the countertop”) while using images as observations of the world — see Figure 1. The direct application of LLMs and VLMs for solving instruction-following tasks remains far from trivial. In particular, the abstract knowledge embedded in LLMs/VLMs often lacks the experiential foundation needed to develop common sense and a grounded understanding of the world [10, 1]. This grounding capacity requires additional skills, such as learning objects’ physical properties and affordances, understanding possible interactions between them, and grasping the temporal causality of the world.
For text-based instruction-based planning, attempts have been made to enhance the relevance of LLMs actions, including prompting strategies such as In-Context Learning (ICL) and Chain-of-Thought (CoT) reasoning [19, 32], as well as the integration of dedicated grounding modules [6, 12, 14]. More recently, reinforcement learning (RL) has also emerged as a powerful method that enables models to acquire real-world knowledge through interaction with the environment [2, 3, 23]. When turning to visual-based instruction-based planning, recent works have tried to apply similar techniques to VLMs [33, 29] by engaging in realistic multimodal environments and fine-tuning using RL. However, the results obtained in [33] remain significantly lower compared to those achieved with text-based models, while [29] relies on detailed text-based descriptions of each image, including object IDs, to guide the VLM during training—a cumbersome and unrealistic requirement in interactive environments.
This paper proposes a method for visual-based instruction-based planning that only requires images as world’s observation, during both training and inference. We introduce the VIPER 111code can be found here framework, which integrate the Visual Perception capabilities of VLMs with the Explainable text Reasoning of LLMs.
VIPER follows a modular approach, as illustrated in Figure 1: (1) a perception module, which generates textual goal-agnostic descriptions from image observations using a frozen VLM, and (2) a reasoning module, which predicts actions based on these descriptions, the task goal, and past interactions. The LLM acts as a policy model, deriving an action distribution from its token probabilities. This structured approach enables explainable, multimodal decision-making, bridging perception and reasoning through text for instruction-based planning.
In summary, our key contributions are as follows:
VIPER is the first method to leverage text as an intermediate representation between perception and reasoning in visual instruction-based planning. We integrate VLM-based perception with LLM-based reasoning, where the VLM describes the image observation to the LLM reasoner, which then generates the appropriate action based on a given textual goal.
We fine-tune the LLM-based reasoning module using a combination of behavioral cloning (BC) with expert trajectories and RL from environment interactions. By leveraging a zero-shot VLM-based perception module to caption image observation, we enhance the decision-making capabilities of the reasoning module without the need for textual supervision from the environment.
Our intermediate text representation enables the development of rich explainability techniques. By detecting important tokens in the image description and the corresponding image regions, VIPER can perform weakly supervised instance segmentation, facilitate a fine-grained analysis of the decision-making process, and provide tools for analyzing failure modes.
Extensive experiments conducted on the challenging ALFWorld environment [20] reveal that VIPER significantly outperforms state-of-the-art visual instruction-based planning methods while narrowing the gap with respect to LLM-based oracles that use precise text descriptions as input. The relevance of our intermediate text representation for explainability is further validated by various quantitative results, while ablation studies validate the importance of our architectural and training choices. Source code will be released upon acceptance.
@@ 2.1 LLM and VLM for Sequential Decision-Making
When tackling long-horizon multimodal sequential decision-making tasks, embodied agents leverage key internal mechanisms such as perception and reasoning to achieve their objectives. Recently, LLMs and VLMs have demonstrated emergent capabilities in understanding complex situations, establishing them as essential architectures for embodied agents [1, 19, 30]. However, these models face challenges related to grounding and alignment with the physical world [16]. A primary line of research focuses on leveraging LLMs for sequential decision-making through text, either by filtering irrelevant actions for a given task [30, 19] or by learning from environmental feedback via reinforcement learning [3, 23]. These methods have demonstrated impressive performance and generalization capabilities. However, they are primarily designed for textual environments, where all state information is transmitted by the environment through text, limiting their applicability to real-world embodied agents. Recently, research has increasingly focused on leveraging VLMs for sequential decision-making tasks in visual environments. For instance, EMMA [29] trained a VLM-based agent for sequential decision-making in visual environments by distilling the actions of a (SOTA) textual agents, such as Reflexion [19], within a parallel textual world. This approach represents an initial step toward utilizing multimodality for instruction generation. However, compared to our work, it remains dependent on a parallel textual environment for task learning, limiting its applicability to real-world scenarios. Closer to our work, RL4VLM [33] employs reinforcement learning to train VLM-based agents within visual environments, enabling them to learn to achieve tasks across diverse environments. However, its performance remains inferior to methods that rely solely on textual observation for reasoning and planning. To address these limitations, we propose a framework that leverages the reasoning capabilities of LLMs in the textual modality and applies them to real-world scenarios through the vision modality. Our approach VIPER creates a decomposition between perception and reasoning via an intermediate text modality: a VLM component processes visual input to generate textual descriptions of all elements present in a scene, and a fine-tuned LLM utilizes this description to generate actions that achieve the specified goal.
@@ 2.2 Text as Intermediate Modality
Recently, VLMs have demonstrated impressive performance on mainstream computer vision tasks such as object detection, visual question answering (VQA), and image captioning [7, 34]. These tasks can be addressed using a single architecture and prompting, making VLMs a promising candidate for robotic applications. However, recent studies have shown that these architectures still underperform compared to their LLM counterparts on purely textual data. [24] demonstrated that VLMs often exhibit inferior reasoning capabilities compared to LLMs, highlighting the advantages of text-based reasoning. For example, [37] utilized a VLM for image captioning before performing a question-answering task to enhance image-related information and improve VQA performance, demonstrating the role of VLMs in extracting scene information. Similarly, [9] employed a VLM to extract bounding boxes and object information from a scene for robotic tasks, while [5] proposed Image-to-PDDL, where a VLM extracts a representation in Planning Domain Description Language (PDDL) [17] of the scene, which is then processed by an LLM to generate robotic actions, effectively leveraging the strengths of both architectures.
@@ 3.3 LLMs as Decision-Making Policies
We leverage the reasoning capabilities of LLMs to solve sequential decision-making tasks [3, 19, 23]. We follow the framework introduced in recent literature [3, 23] to derive a policy (a probability distribution over possible actions) from an LLM, as shown in Figure 2(c). More precisely, given a prompt , the probability of an action sequence , denoted as , is obtained from the decoding probabilities of the tokens composing . Specifically, this probability is computed as:
@@ 3.4 Training Strategies
Recent works [3, 2] show that while LLMs possess common-sense knowledge, they lack grounding and are misaligned with physical environments [16], leading to poor zero-shot performance in complex interactive decision-making tasks. To address this, following [3, 23], we fine-tune the reasoning module (LLM) using a sequential strategy. First, we align it with a good reference policy using BC on a dataset of expert demonstrations, which is assumed to be available in the environment. Then, we refine its performance through RL via exploration and interaction with the environment. This BC+RL approach is well-established in the control community [15] and has demonstrated its relevance in our ablation study (see Section 4.3). While VIPER’s reasoning module operates on textual inputs, the textual description of the image observation is provided by a frozen VLM. Therefore, unlike EMMA [29], which trains its model by distilling a text expert policy and thus requires textual observations from the environment, our training relies solely on the goal and the image observations .
BC fine-tuning: We pre-train the LLM policy parametrized by a parameter , using a dataset of expert demonstrations of observation-action trajectories through BC. In our case, this expert dataset is created by selecting the successful trajectories of a rule-based expert (Sup. 6). The BC objective is to maximize the log-likelihood of expert actions with respect to the action distribution predicted by the LLM policy along the expert trajectory. Therefore, the BC loss is the negative log-likelihood, defined as:
@@ 3.5 Explaining Agent Decision
Furthermore, we perform weakly-supervised instance segmentation of important concepts in the image description used by the reasoning module by computing the integrated gradient [22] of the chosen action with respect to each text token in the image description, and selecting the most significant ones. Following the same procedure, we compute the integrated gradient of the selected tokens, with respect to the image (see Sup. D for more details). This provides a rich tool for explainability, linking the performance of the perception and reasoning modules, and enabling to finely analyze failure cases (see Sec. 4.4).
@@ 4 Experimental Results
| Environnement | Pick | Look | Clean | Heat | Cool | Pick2 | Avg | |
| AutoGen* [27] | Text | 0.92 (-) | 0.83 (-) | 0.74 (-) | 0.78 (-) | 0.86 (-) | 0.41 (-) | 0.77 (-) |
| ReAct* [30] | Text | 0.71 (18.1) | 0.28 (23.7) | 0.65 (18.8) | 0.62 (18.2) | 0.44 (23.2) | 0.35 (25.5) | 0.54 (20.6) |
| DEPS* [25] | Text | 0.93 (-) | 1.00 (-) | 0.50 (-) | 0.80 (-) | 1.00 (-) | 0.00 (-) | 0.76 (-) |
| Reflexion* [19] | Text | 0.96 (17.4) | 0.94 (16.9) | 1.00 (17.0) | 0.81 (19.4) | 0.83 (21.6) | 0.88 (21.6) | 0.91 (18.7) |
| EMMA* [29] | Text + Vision | 0.71 (19.3) | 0.88 (19.6) | 0.94 (17.5) | 0.85 (19.6) | 0.83 (19.9) | 0.67 (22.4) | 0.82 (19.5) |
| Florence-2 [28] | Vision | 0.00 (30.0) | 0.06 (28.5) | 0.0 (30.0) | 0.0 (30.0) | 0.0 (30.0) | 0.0 (30.0) | 0.01 (29.7) |
| Idefics-2 [13] | Vision | 0.04 (29.2) | 0.06 (28.2) | 0.0 (30.0) | 0.0 (30.0) | 0.0 (30.0) | 0.0 (30.0) | 0.02 (29.5) |
| MiniGPT-4* [36] | Vision | 0.04 (29.0) | 0.17 (17.7) | 0.0 (30.0) | 0.19 (26.3) | 0.17 (26.7) | 0.06(28.9) | 0.16 (26.9) |
| InstructBLIP* [4] | Vision | 0.50 (21.5) | 0.17 (26.8) | 0.26 (25.0) | 0.23 (27.2) | 0.06 (28.9) | 0 (30.0) | 0.22 (26.2) |
| RL4VLM* [33] | Vision | 0.47 (-) | 0.14 (-) | 0.10 (-) | 0.14 (-) | 0.18 (-) | 0.18 (-) | 0.21 (-) |
| VIPER | Vision | 0.80 (13.1) | 0.77 (16.7) | 0.77 (19.5) | 0.92 (14.3) | 0.71 (20.4) | 0.53 (24.0) | 0.75 (18.0) |
@@ 4.1 Experimental Setup
Environment. Our experiments are conducted on the ALFWorld benchmark [20], a 3D environment that simulates six categories of household tasks with textual goals. ALFWorld has become a standard benchmark in the literature for evaluating models in long-horizon multimodal decision-making tasks, offering a realistic and interactive environment well suited for studying real-world embodied applications. This benchmark contains sets of training, in-distribution validation, and out-of-distribution (OOD) episodes that encapsulate specific layouts of the environment (room in the house, object and furniture placement, etc.). We follow the standard protocol and train VIPER on the training set of episodes, perform cross-validation on the in-distribution validation set, and report our results on the OOD test set. We evaluate the models using the mean success rate (higher is better, ) and the mean length of successful trajectories (lower is better, ). See Sup. 6 for more details about the environment.
Architecture and Training. VIPER uses Florence-2-220M [28] as VLM and Mistral-7B [11] as LLM, with further discussions on model choices in Section 4.3. We fine-tune the LLM as described in Section 3.4 using low-rank adaptation (LoRA) with parameters and on the attention layers. For BC, we train the model until convergence and for PPO for approximately interaction steps with the environment. Both fine-tuning processes use a batch size of 2 and the Adam optimizer with a learning rate of . For each task, BC fine-tuning takes and PPO on 4 H100 GPUs.
Baselines. We compare VIPER with various SOTA models from the literature, categorized into two groups. The first group comprises models trained with supervision from a parallel textual environment, including Reflexion [19], DEPS [25], ReAct [30], AutoGen [27] which only receive textual observation of the environment and EMMA [29] which uses both visual and textual observation during training but only access visual observation during inference. Like VIPER, the second group consists of models that exclusively rely on visual information from the environment, including zero-shot models (MiniGPT-4 [36], InstructBLIP [4], Florence 2 [28], Idefics 2 [13]) and a VLM fine-tuned through BC and RL, RL4VLM [33]. All methods of the first group accessing the textual observations benefit from richer scene information, including the number of objects, object IDs within the description, and the correspondence between object IDs and actions. In contrast, the models of the second group, including the zero-shot perception module in VIPER, cannot perfectly generate this level of detail, as this information is internal to the environment and not directly observable from images. Therefore, the models of the first group can be seen as oracle baselines.
@@ 4.2 Comparison with the state-of-the-art
We compare VIPER with the two baseline families on ALFWorld (see Table 1), evaluating both the mean success rate and the mean trajectory length required to achieve the goals. As observed, VIPER outperforms zero-shot visual-only methods, achieving an average 50pt improvement in success rate and an 8-step reduction in trajectory length across multiple tasks. It also surpasses RL4VLM [33], the SOTA method most closely related to our work that applies both BC and RL fine-tuning. This lower performance of RL4VLM [33] may be explained by the use of CoT reasoning with free generation for action prediction, which prevents it from being constrained to the action space. This leads to irrelevant predictions, as explained in Section 3.3. Another factor could be the absence of past action history in the agent’s prompt, as discussed in Sup. B.2. Besides, VIPER outperforms ReAct [30] and matches the performance of AutoGen [27] and DEPS [25], which rely on textual observations from the environment. Moreover, VIPER approaches the performance of oracle methods such as EMMA [29] and Reflexion [19], achieving more than 70% success rate in the first five tasks and more than 50% in the Pick2 task. The performance gap between Pick2 and the other tasks results from the increased complexity of manipulating multiple objects. Text-based methods, such as EMMA and Reflexion, do not face this challenge, as they have access
@@
Comparisons between textual observations and VIPER-generated descriptions can be found in Sup. 6. Furthermore, VIPER achieves the shortest mean episode length compared to both the oracle and vision-only baselines, demonstrating that its successful trajectories are more efficient, taking more goal-directed actions and avoiding unnecessary steps.
(a) and incorrect actions
(b). VIPER’s intermediate text modality filters relevant elements for action generation and retrieves them from the image using integrated gradients, helping to analyze both correct (e.g., identifying cups and sinks in (a)) and incorrect predictions (e.g., missing the egg while focusing on the irrelevant apple in (b)).4.3 Ablation study
Impact of BC and RL Training. Through an ablation study, we evaluate the impact of each training strategy defined in Section 3.4 – BC and RL – by comparing different variants of VIPER in Figure 3. Specifically, we report the performance of VIPER trained only with BC (VIPER BC) using rule-based expert demonstrations (see Sup. 6), only PPO (VIPER PPO), and the sequential approach where PPO follows BC (VIPER BC+PPO). Although using SOTA VLMs and LLMs, the zero-shot model performs poorly across all tasks, highlighting the complexity of the tasks and the necessity of fine-tuning. The results further show that while applying RL fine-tuning alone (VIPER PPO) is sufficient to learn simpler tasks (Pick and Look), using BC beforehand is essential to learn more complex ones (Clean, Heat Cool and Pick2). Furthermore, while VIPER BC already outperforms vision-only SOTA models, adding RL further improves performance, producing a boost of 3pt to 9pt depending on the task. This performance gain is larger when considering weaker reasoning modules (see Sup. C).